OpenStudy (bbb911):

This is geometry: The length of the hypotenuse of an isosceles right triangle is 30 meters. Find the area of the triangle. Round to the nearest tenth, if necessary.

OpenStudy (anonymous):

Area of the triangle : A=(1/2)*base* height in an right isosceles triangle, the base is one leg and the height is the other leg

OpenStudy (anonymous):

A=(1/2)leg^2= (1/2)*(30/sqrt(2))^2=30^2/4

OpenStudy (anonymous):

all you have to do is solve.

OpenStudy (anonymous):

so just solve 30^2/4

OpenStudy (anonymous):

and since the legs are the same you know \[a^2+a^2=30^2\] \[2a^2=30^2\] \[a=\frac{30}{\sqrt{2}}\]

OpenStudy (anonymous):

those are the lengths of the legs multiply them together and divide by two to get \[\frac{1}{2}\times\frac{30^2}{2}=\frac{900}{4}=225\]